The problem is:
"Find 2 consecutive integers such that the smaller plus twice the larger equals -61."
If the smaller integer equals S, this is what I have come up with:
S + (s+1x2) Is this correct?
I am not sure if the consecutive number is to the power of 2 or the consecutive number +1 is to the power of 2… can you advise?
Thank you!!!

Answer by richard1234(7193)

The next number (the larger one) is the one that is being multiplied (multiplied by 2, not raised to the power of 2). The equation should look like

S + 2(S+1) = -61

3S + 2 = -61

3S = -63

S = -21, S+1 = -20.